Does anyone know how to do this problem
Bob has just finished climbing a sheer cliff above
Answer #1

81 mph = 118.8 ft/s

time of flight = 2vsintheta/g

0.910 = 2*118.8 * sin theta /32

theta = 7.04 degree

Vx = V cos theta = 117.9 ft/s

Vy = 14.56 ft/s

since the range is 387 ft

t = R/Vx = 3.28s

total time from cliff to ground = 3.28 - 0.910 = 2.37 s

s = vy*t + 1/2*gt^2

s = 124.38 ft